Database last updated on Sun, 01 Feb 2026 22:40:50 +0000 by localhost

Instrument constant changes for the week before 2023238 (3 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 29 rows)

Measurements and tests summary.


_jday_loc_serial_skcedfmhghpsl
238AMUNDSEN-SCOTT#021Next schedule will be spmoon (now spmoon ZS= 0 ZS= 100.342),Next schedule will be spmoon (now spmoon ZS= 0 ZS= 100.401),Next schedule will be spmoon (now spmoon ZS= 0 ZS= 100.461),spmoon21126616
(0 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (1 sec)

File nameCountSerialExpected (+/-20) Measured 30-day history
B23823.02116#021596578